Calculate the number of mmoles of limonene used in the reaction (1.3 g limonene used).
Answer #1

limonene formula is C10H16

Molar mass of C10H16,

MM = 10*MM(C) + 16*MM(H)

= 10*12.01 + 16*1.008

= 136.228 g/mol

mass(C10H16)= 1.3 g

use:

number of mol of C10H16,

n = mass of C10H16/molar mass of C10H16

=(1.3 g)/(1.362*10^2 g/mol)

= 9.543*10^-3 mol

= 9.543 mmol

Answer: 9.5 mmol